Back to the NIKA2 main page

Daily Reports

Oct 10th, Monday

Special session before the beginning of the NIKA2 Pool - Summer 2022

10:30-13:00; Martino, Alessandro, Dave, Santiago, Angel: (transcription of the e-mail from Martino) "(...)this morning we have prepared the new sweeps for NIKA2. They are ready to be used: when starting the acquisition with the config file 'run73.ini' you'll get the 'new sweep'. To use the 'old sweep', you have to use the 'run72.ini' configuration file. Unfortunately, it looks like tonight the opacity could be pretty high, but let's hope for the best. Just as a reminder: the new sweep is made with the cryostat window closed."

Before the official time for starting the scheduled tests of the Pool, we started the DAQ twice to get some samples of noise traces (NIKA2 window still being closed, telescope stopped due to bad weather) using the run72.ini ('old sweep') and run73.ini ('new sweep') configuration files. Corresponding screen-shots are given below.

13/10/22; In order to simplify swapping procedures between old and new sweep or viceversa, and keep the same run number after the last cooling down (72), new filenames have been defined in the nika2@nika2-a:~/NIKA/Processing/NIKA_lib_V3_IRAM/Acquisition/experiments/ folder:

* To activate the old sweep, total power, do 'cp -p run72_tp_old_sweep.ini run72.ini'

* To activate the new sweep, total power, do 'cp -p run72_tp_new_sweep.ini run72.ini'

Change 'tp' to 'pol' in the previous commands for polarization.

Monitor status and noise traces, 'old sweep', bad KIDs Ar1: 36/1055; Ar2: 172/573; Ar3: 39/1056.

attachment:traces_run72_Ar1_20221010.png attachment:traces_run72_Ar2_20221010.png attachment:traces_run72_Ar3_20221010.png

Monitor status and noise traces, 'new sweep', bad KIDs Ar1: 31/1061; Ar2: 7/581; Ar3: 57/1077.

attachment:traces_run73_Ar1_20221010.png attachment:traces_run73_Ar2_20221010.png attachment:traces_run73_Ar3_20221010.png

18:00-22:00; Angel: tonight's plan includes a discrete pointing session, getting beam maps and other calibrations with NIKA2. If conditions are cool, we would start science observations. Unfortunately, the telescope is stopped since 13:30 (rainy, foggy). At 18:00, tau_225GHz > 1 with thick cloud intervals. Conditions improved (tau_225GHz ~0.6) at ~19h for EMIR-backup project 019-22 (3 mm). Forecasts improved a bit too. G34.3+0.2 line calibrator, reference off-position, and EMIR observed once (submap5_beta) for project 019-22. The target is too low for continuing. Switched to project 083-22. Opacity (225 GHz) ~0.7 and a bit unstable. Scans 20221010s174 and s175 got stuck. After pointing/focus, 3 reps. on 39568-high were observed.

22:30-03:00; Pablo: Opacity dropped to ~0.6; EMIR backup project 018-22 (3 mm, Psw) observed. One rep. of domapC1 and domapC2, both with setup 1. Telescope got stuck when OTFmap 20221010s222 was running. Repeated. Switching to project 055-22. Onoff-00 calibration on L183 after pointing/focus on Mars, according to the expected. Script l183otf 290 -230 launched twice.

Oct 11th, Tuesday

03:00-07:00; Angel: Continuing on EMIR backup project 055-22; tau_225GHz ~0.9; tauZenit at ~100 GHz (MRTCAL) ~0.2; very stable atmosphere. Two reps. of script l183otf 290 -360 on L183 (provided maps completed, TBC). Switching to project 051-22. Nice pointing/focus on Mars (spindle vertical failure -no tracking- before getting the focus correction: s28-s30); tauZenit at ~100 GHz (MRTCAL) < 0.2. Scan 20221011s40 got stuck. NGC 1977 observed: 05-otf-NGC1977_E090-map[1,2,3]-hori (1 rep. each), 05-otf-NGC1977_E090-map[1,2,3]-ver (1 rep. each); After 07h, tauZenit at ~100 GHz (MRTCAL) > 0.4; tau_225GHz > 1.4. Stopping at 07:20.

DailyReportsNika2Pool101022 (last edited 2022-10-19 11:06:45 by NikaBolometer)